OpenStudy (anonymous):

Find the constant of variation for the relation, and use it to write an equation for the statement. Then solve the equation. If y varies directly as x, and y = 6 when x = 7, find y when x=5.

OpenStudy (jdoe0001):

something varies directly to something else, usually means \(\bf \textit{something} = n \times \textit{something else} \implies y = n \times x\) so, what's the "constant of variation" \(\bf \text{n}\) ? we dunno but we do know that, when x = 7, y = 6 so let's use that in our equation \(\bf y = n \times x \implies 6 = n \times 7\) solve for "n", to find the \(\bf \text{contant}\) of variation

OpenStudy (jdoe0001):

ahem \(\bf \text{constant}\) rather =)

OpenStudy (anonymous):

so n=6/7

OpenStudy (anonymous):

what is a statement of a triangle inequality theorem?

OpenStudy (jdoe0001):

yes, so now that you know that, use that in the equation, to find "y" when x = 5 so, x = 5, what's "y"?, well \(\bf y = n \times x \implies y = \cfrac{6}{7}\times 5\)

OpenStudy (anonymous):

Solve y = kx for k when y=6 and x=7.\[y=\frac{6 x}{7} \] find y when x=5

OpenStudy (anonymous):

y=30/7

OpenStudy (anonymous):

thank you guys!

OpenStudy (anonymous):

np